if a body which is uniformly retarded comes to rest in five secs after travelling a distance of 10m. what is its initial velocity?
1
Expert's answer
2018-12-17T07:43:10-0500

The distance body traveled

s=(v_i+v_f)/2 t

Thus

v_i=2s/t-v_f=(2×10m)/5s-0 m/s=4 m/s